# Premerger phenomena in neutron-star binary coalescences

### v2

@@ IV The mechanics of late inspirals: gravitational waves
Consider a binary, involving at least one neutron star, with component masses $M_{A}$ and $M_{B}$. Ignoring complications about how the objects may reach short orbital separations $a$ [this issue is reminiscent of the famous “final parsec problem”, though for neutron-star mergers the solution is likely rooted in common-envelope theory; 393], the (quadrupole formula) GW inspiral time reads
@@
Since the pioneering works of Peters and Mathews 310, Peters 309, it is anticipated that compact binary-mergers circularise well before merger, so that the eccentricity $e\ll 1$. Note the normalization in expression (11): although it would only take light $\sim 1$ second to travel between two such stars, GW radiation-reaction takes $\approx 10$ Myr. Narrowing our attention immediately to late stages where the separation is $\lesssim 20R_{A}$ for “canonical values” of the stellar radius $R_{A}=12$ km and mass $M_{A}=1.4M_{\odot}$ (see Sec. II.1), we find $\tau_{\rm GW}$ from (11) corresponds to tens of seconds, or more precisely that
@@
The convergence to coalescence, occurring roughly when $a\lesssim 3(R_{A}+R_{B})/2$ [e.g. 237, 236], accelerates rapidly in the final stages. The orbital separation can be inversely related to the orbital frequency $\Omega_{\rm orb}$ through a Keplerian or quasi-Keplerian relationship, and thus $\Omega_{\rm orb}$ grows in time. This sweep-up behaviour culminates in what is known as the “chirp” in the GW community.
As we explore throughout the remainder of this Section, the simple result (12) is an overestimate [e.g. 214]. When considering PN or finite-size effects pertaining to tides, the inspiral time is reduced as additional means of energy depletion become available: the impact of tides increases as the orbital frequency sweeps up because the tidal forcing terms similarly increase in frequency. Getting a precise estimate of the tidal dephasing is of critical importance when trying to connect to observational data, since without accurate inference of the time relative to merger (in some appropriate frame of reference) that some high-energy events occurs, one cannot hope to extract the maximum amount of information. More generally though, the tides encapsulate details about stellar structure and thus matching dephasing templates to data can be used to learn about many areas of fundamental physics.
@@ IV.1 Tides: general theory
Within the literature, one can find a few different definitions for the qualifiers “adiabatic”, equilibrium”, and “dynamical” in the context of tides. For example, following Yu et al. 441, the word “adiabatic” emphasises that one considers deformations in the zero-frequency limit $\Omega_{\rm orb}\rightarrow 0$. This is more typically called the equilibrium tide, though Yu et al. 441 make the distinction by allowing the latter to include $\Omega_{\rm orb}>0$ influences. Finite-frequency terms involve both non-resonant and resonant excitations, the latter of which these authors reserve for “dynamical” tide status. In other works, the inclusion of non-resonant, but time-dependent terms leads to corrections of the equilibrium quality factors which are then called “effective” [e.g. 314]. In this Review, the word “dynamical” conveys $\Omega_{\rm orb}>0$, though we are mostly interested in resonances and their connections to high-energy phenomena. We emphasise at this stage that many of the formulae that we present are done so in a Newtonian language. When one reaches the level of wishing to actually compare results with data in a serious way, it will generally be necessary to consider fully GR expressions insofar as possible [^4] Subtle issues arise in GR, making this non-trivial. For example, even the notion of eccentricity is ambiguous [213]. An all-encompassing notion of “tide” is also hard to define via the Weyl tensor or other means; see Section I.C in Pitre and Poisson 313.. Given that our main purpose here is pedagogical, we avoid writing out formalisms in this way, though point the interested reader toward relevant literature where appropriate. A depiction of these two class of tides is given in Figure 6.
@@
Following Lai and Wu 238 and others, we orient a spherical coordinate system around (the centre of mass of) $M_{A}$ with the “$z$-axis” directed along L. The gravitational potential produced by $M_{B}$ can then be expanded in terms of spherical harmonics $Y_{\ell m}$ [e.g. 321, 333]
@@
where $\mathcal{D}^{(\ell)}_{mm^{\prime}}$ is the so-called Wigner $\mathcal{D}$-function [e.g. 190, 434, 230], and $\phi_{s}=\phi+|\boldsymbol{\Omega}_{s}|t$.
@@ IV.1.1 Mathematical description and calculation methods
Modes with $n>1$ are referred to as “overtones”, as $n$ is defined by counting the number of radial nodes in the eigenfunction [see, e.g. 409]. Rotation, however, generally makes a separation of variables as above impossible, and a further sum over a dummy azimuthal index becomes necessary unless one further sets up a hierarchy in powers of $|\boldsymbol{\Omega}_{s}|$ [e.g. 370].
@@
In solving mode problems one must also take care to note that complex conjugates generally also solve the master equation, though we ignore this complication here for pedagogical purposes. [see 314, for a recent discussion].
At this stage, the amplitude of the modes have not yet entered, as these fall out of the homogeneous equation (18). Enter the tides. Formally, their accounting amounts to instead solving the inhomogenous version of the master equation,
@@
Fortunately, except possibly at very late stages in the inspiral, the tidal distortion of the spectrum is weak [114], as can be formally estimated with the Unno et al. 409 formula; see Sec. IV.2. One thus considers a “volume-averaged” problem where the amplitude evolution, $q_{\alpha}(t)$, of each mode is considered in isolation. The problem is thus reduced from 1+3 to 1+0 dimensions, and we will end up with some ODEs for the amplitude evolution. The key step involves introducing an inner product,
@@
where the overlap integrals [^6] Two ways of generalising overlaps to GR have been proposed in Kuan et al. 230 and Miao et al. 273. The orthogonality between modes and the sum rule for tidal overlaps [332] are equally respected by both definitions, while only the former predicts a vanishing $g$-mode overlap in the zero stratification limit (at least for a simple, constant $\delta$ law; see Sec. III.3). This issue is related to overlap “leakage” discussed in Sec. VI.4.1. Throughout we adopt the definition of Kuan et al. 230. read
@@
Table 2: A non-exhaustive list of Post-Newtonian effects that influence the GW inspiral for compact objects. Constraints on each of these orders can be found in the violin plots produced by the LVK consortium; see Abbott et al. 8. In principle, stellar-structure and multipole moments corrections and additional couplings enter at ever-higher orders [401].
@@
0 Energy deposited into modes Alexander 16, 17
0.5 Scalar contributions to dynamics (non-GR) Yagi et al. 436
1 Stellar structure; pericentre advance Wagoner and Malone 421, Andersson et al. 28
1.5 Spin-orbit coupling; tail backscatter Blanchet et al. 66
2 Self-spin, spin-spin, quadrupole-monopole couplings, mag. dipoles Poisson 315, Ioka and Taniguchi 197
3 Gravitational tails of tails Blanchet 64
4 Dissipative tidal number Ripley et al. 337, Saketh et al. 344
5 Gravitoelectric quadrupole Love number Damour et al. 107, Yagi 435
$6$ Gravitomagnetic quadrupole Love number Abdelsalhin et al. 10, Henry et al. 181
$6.5$ Spin-tidal coupling Abdelsalhin et al. 10, Castro et al. 78
Thus, provided the free mode spectrum can be constructed, one need “only” solve equation (27). This, however, is still not quite the end of the story as the orbital phase $\Phi$ must be evolved simultaneously. By modelling the fluid motion in the neutron star as a set of harmonic oscillators [16, 17, 347] and incorporating mode kinetics into the Hamiltonian of the binary, Schäfer 346, Kokkotas and Schafer 217, Kuan et al. 230 have shown how this can be achieved with a high-order PN method, though with the modes themselves calculated in full GR without Cowling (i.e. including the metric perturbations). While not all are taken into account [such as those occurring at negative orders in gravities with non-metric fields where dominant mono- or dipolar radiation exists; 436], we list PN orders at which various effects occur in Table 2. We remark in this respect that although, strictly speaking, radiation-reaction is a 2.5 PN effect, often this is said to appear at Newtonian order when using a Newtonian definition for the system’s quadrupole moment(s) (see below). There is some ambiguity therefore in what constitutes a “PN order” exactly.
PN effects in shaping stellar structure have been examined in several references [e.g. 421, 374, 28] and starts already at first order, though no systematic analysis of how these directly influence the inspiral has been carried out. Tides and spins contribute hierarchically at several PN orders accounting, for instance, for the breakdown of the point-particle approximation. Tides absorb orbital energy in the Newtonian theory already (i.e. at $0^{\rm th}$ order), while spin couples to the orbital angular momentum at the 1.5th order and to self-spin and companion spin at 2nd order. At 2.5th order enters the leading-order dissipation due to GW emission (cf. equation 11). The electric- (magnetic-) type stationary deformations factor into the dynamics at 5th (6th) order, as introduced in Sec. IV.1.2, and the coupling between this deformation to the spin follows at the next half order. GW170817 placed constraints on each of these PN orders, as detailed in Abbott et al. 8.
With the above in mind, the notion of dephasing can be made precise. In the stationary phase approximation, the frequency-space GW phase, $\Psi$, can be written as [e.g. 142, 102]
@@ IV.1.2 Tides: equilibrium
The equilibrium tide simply corresponds to the “$\Phi(t)=0$” portion of the dynamics. In this case, the quasiharmonic time dependence of (13) drops out, and our interest shifts to bulk, geometric deformations of the stellar surface (“zero-frequency oscillations”). The extent to which the stellar interior is suspectible to a time-independent tidal potential can be encapsulated by the (shape) Love numbers; effectively, much like the $Q_{\alpha}$ defined previously, these measure the extent of orthogonality between the stellar fluid (or solid) and some angular portion of the tidal field (i.e. $\ell,m^{\prime}$). In general, therefore, there is an infinite set of Love numbers, though often in the literature one will find the term “Love number” just to mean the quadrupolar, $m^{\prime}=0$ Love number. In the context of compact binaries in GR, Damour 104 first quantified how tidal Love numbers influence inspiral. Note also that if the star is static and axisymmetric at background order, the index $m^{\prime}$ falls out of the equations and thus the static, or even effective, Love numbers are characterised by a single harmonic number $\ell$. These reduced Love numbers are typically denoted $k_{\ell}$.
@@
which is the same $U$ from (13). Note that the multipole moments $I_{\ell m}$ are also implicitly defined at some PN order; see Section V in Thorne 401. At Newtonian order, $I_{\ell m}$ is defined through an integral over the mass-density weighted by a spherical harmonic and radius to power $\ell$:
@@
where the (axisymmetric) mass ($M_{\ell}$) and current ($S_{\ell})$ multipole moments can be defined either via the Thorne 401 or Geroch-Hansen [178] formulae. These formulations are totally equivalent by a theorem due to Gürsel 174. In the above we have the odd and even parity sectors of contractions of the trace-free Weyl tensor [e.g. 108, 297]. Using the schemes introduced by Pappas and Sotiriou 299, Suvorov and Melatos 388 to generalise the Geroch-Hansen definitions to theories beyond GR, one could try to extend the correspondence (38) to some other theory of gravity. This has not yet been attempted.
Figure 7: The main radial gravitational potential $A(R)$ — the lapse function of the effective-one-body spacetime over which the dynamics of a test particle describes the dynamics of the real binary system [71, 72] — as a function of the radial coordinate $R=a$ with $M=M_{A}+M_{B}$ (for this Figure only) in a tidal-EOB model. The binary black hole sector of this model (solid black) shows a deviation from the lapse of the Schwarzschild spacetime (dotted), indicating the dynamics of a test particle on the effective spacetime is different from that of a (plunging) orbit around a Schwarzschild black hole due to the finite symmetric mass ratio $\nu=M_{A}M_{B}/M^{2}$ of the binary (which should not be confused with the our use of $\nu$ as stellar spin used elsewhere). For the neutron star case, the gravitational potential including up to next-to-next-to-leading-order tidal effects (i.e. up to 7 PN; green dashed) is shown together with the result resumming the gravitational-self-force information up to 7.5 PN [61, red]. The specific equal-mass neutron-star binary considered here has a tidal deformability (equation 39) of $\kappa_{2}^{T}\approx 73.55$. From Bernuzzi et al. 55 with permission.
Tides effectively enhance the “attraction” between the components of the binary, leading to earlier merger relative to binary black-holes or point-particles. The tidal dynamics are elegantly described by a radial interaction potential in the effective-one-body (EOB) formalism [62, 55, 187]. PN models of tides in the EOB formalism were initially verified against numerical simulations by Baiotti et al. 43, Bernuzzi et al. 56, who demonstrated that the models can be improved with such calibrations but that it is not possible to accurately describe tides close to merger with a PN model. Such a restriction was lifted by Bernuzzi et al. 55 through resumming techniques. A schematic depiction of this attractive feature is shown in Figure 7, where the red curve describes the radial potential for a neutron-star binary with tidal resumming. A significant drop in the potential occurs at larger radius, indicating the stronger attraction felt by the binary than the equivalent binary black hole. The dominant term of the tidal imprints on waveform is delivered by the quadrupole quantity [107]
@@
for $\Lambda_{A}=\frac{2}{3}k_{2}^{A}(R_{A}/M_{A})^{5}$ with the same definition for object $B$ [see also 185, 143, 186, 106]. Note $k_{2}$ is defined in expression (34), effectively being the constant of proportionality between the quadrupole moment and the quadrupolar tidal potential (appropriately generalised to GR via the Weyl tensor; though cf. Footnote 4).

### v3

@@ IV The mechanics of late inspirals: gravitational waves
Consider a binary, involving at least one neutron star, with component masses $M_{A}$ and $M_{B}$. Ignoring complications about how the objects may reach short orbital separations $a$ [this issue is reminiscent of the famous “final parsec problem”, though for neutron-star mergers the solution is likely rooted in common-envelope theory; 407], the (quadrupole formula) GW inspiral time reads
@@
Since the pioneering works of Peters and Mathews 323, Peters 322, it is anticipated that compact binary-mergers circularise well before merger, so that the eccentricity $e\ll 1$. Note the normalization in expression (11): although it would only take light $\sim 1$ second to travel between two such stars, GW radiation-reaction takes $\approx 10$ Myr. Narrowing our attention immediately to late stages where the separation is $\lesssim 20R_{A}$ for “canonical values” of the stellar radius $R_{A}=12$ km and mass $M_{A}=1.4M_{\odot}$ (see Sec. II.1), we find $\tau_{\rm GW}$ from (11) corresponds to tens of seconds, or more precisely that
@@
The convergence to coalescence, occurring roughly when $a\lesssim 3(R_{A}+R_{B})/2$ [e.g. 246, 245], accelerates rapidly in the final stages. The orbital separation can be inversely related to the orbital frequency $\Omega_{\rm orb}$ through a Keplerian or quasi-Keplerian relationship, and thus $\Omega_{\rm orb}$ grows in time. This sweep-up behaviour culminates in what is known as the “chirp” in the GW community.
As we explore throughout the remainder of this Section, the simple result (12) is an overestimate [e.g. 222]. When considering PN or finite-size effects pertaining to tides, the inspiral time is reduced as additional means of energy depletion become available: the impact of tides increases as the orbital frequency sweeps up because the tidal forcing terms similarly increase in frequency. Getting a precise estimate of the tidal dephasing is of critical importance when trying to connect to observational data, since without accurate inference of the time relative to merger (in some appropriate frame of reference) that some high-energy events occurs, one cannot hope to extract the maximum amount of information. More generally though, the tides encapsulate details about stellar structure and thus matching dephasing templates to data can be used to learn about many areas of fundamental physics.
@@ IV.1 Tides: general theory
Within the literature, one can find a few different definitions for the qualifiers “adiabatic”, equilibrium”, and “dynamical” in the context of tides. For example, following Yu et al. 455, the word “adiabatic” emphasises that one considers deformations in the zero-frequency limit $\Omega_{\rm orb}\rightarrow 0$. This is more typically called the equilibrium tide, though Yu et al. 455 make the distinction by allowing the latter to include $\Omega_{\rm orb}>0$ influences. Finite-frequency terms involve both non-resonant and resonant excitations, the latter of which these authors reserve for “dynamical” tide status. In other works, the inclusion of non-resonant, but time-dependent terms leads to corrections of the equilibrium quality factors which are then called “effective” [e.g. 327]. In this Review, the word “dynamical” conveys $\Omega_{\rm orb}>0$, though we are mostly interested in resonances and their connections to high-energy phenomena. We emphasise at this stage that many of the formulae that we present are done so in a Newtonian language. When one reaches the level of wishing to actually compare results with data in a serious way, it will generally be necessary to consider fully GR expressions insofar as possible [^4] Subtle issues arise in GR, making this non-trivial. For example, even the notion of eccentricity is ambiguous [221]. An all-encompassing notion of “tide” is also hard to define via the Weyl tensor or other means; see Section I.C in Pitre and Poisson 326.. Given that our main purpose here is pedagogical, we avoid writing out formalisms in this way, though point the interested reader toward relevant literature where appropriate. A depiction of these two class of tides is given in Figure 6.
@@
Following Lai and Wu 247 and others, we orient a spherical coordinate system around (the centre of mass of) $M_{A}$ with the “$z$-axis” directed along ${L}$. The gravitational potential produced by $M_{B}$ can then be expanded in terms of spherical harmonics $Y_{\ell m}$ [e.g. 334, 347]
@@
where $\mathcal{D}^{(\ell)}_{mm^{\prime}}$ is the so-called Wigner $\mathcal{D}$-function [e.g. 197, 448, 239], and $\phi_{s}=\phi+|\boldsymbol{\Omega}_{s}|t$.
@@ IV.1.1 Mathematical description and calculation methods
Modes with $n>1$ are referred to as “overtones”, as $n$ is defined by counting the number of radial nodes in the eigenfunction [see, e.g. 423]. Rotation, however, generally makes a separation of variables as above impossible, and a further sum over a dummy azimuthal index becomes necessary unless one further sets up a hierarchy in powers of $|\boldsymbol{\Omega}_{s}|$ [e.g. 384].
@@
In solving mode problems one must also take care to note that complex conjugates generally also solve the master equation, though we ignore this complication here for pedagogical purposes. [see 327, for a recent discussion].
At this stage, the amplitude of the modes have not yet entered, as these fall out of the homogeneous equation (18). Enter the tides. Formally this amounts to instead solving the inhomogenous version of the master equation,
@@
Fortunately, except possibly at very late stages in the inspiral, the tidal distortion of the spectrum is weak [114], as can be formally estimated with the Unno et al. 423 formula; see Sec. IV.2. One thus considers a “volume-averaged” problem where the amplitude evolution, $q_{\alpha}(t)$, of each mode is considered in isolation. The problem is thus reduced from 1+3 to 1+0 dimensions, and we will end up with some ODEs for the amplitude evolution. The key step involves introducing an inner-product,
@@
where the overlap integrals [^6] Two ways of generalising overlaps to GR have been proposed in Kuan et al. 239 and Miao et al. 284. The orthogonality between modes and the sum rule for tidal overlaps [346] are equally respected by both definitions, while only the former predicts a vanishing $g$-mode overlap in the zero stratification limit (at least for a simple, constant $\delta$ law; see Sec. III.3). This issue is related to overlap “leakage” discussed in Sec. VI.4.1. Throughout we adopt the definition of Kuan et al. 239. read
@@
Table 2: A non-exhaustive list of Post-Newtonian effects that influence the inspiral for compact objects. The order counting presented in the first column is quoted with reference to the power of $c$ of the associated coefficient(s) in the equations of motion for binaries (i.e., $N$ PN corresponds to a $c^{-2N}$ scaling). Constraints on each of these orders can be found in the violin plots produced by the LVK consortium; see [8]. In principle, stellar structure and multipole moments corrections together with additional couplings enter at ever-higher orders [415]. Some additional modified gravity terms may also enter at lower PN orders; see Table 3 in Yunes, Yagi, & Pretorius [458].
@@
0 Energy deposited into modes [16, 17]
1 Stellar structure; pericentre advance [435, 28]
1.5 Scalar-field contributions to dynamics (non-GR) [288, 450]
1.5 Spin-orbit coupling; tail backscatter [65]
2 Self-spin, spin-spin, quadrupole-monopole couplings, mag. dipoles [328, 204]
2.5 Quadrupole formula for GWs [323]
3 Gravitational tails of tails [63]
4 Dissipative tidal number [351, 358]
5 Gravitoelectric quadrupole Love number [106, 449]
$6$ Gravitomagnetic quadrupole Love number [10, 187]
$6.5$ Spin-tidal coupling [10, 77]
Thus, provided the free mode spectrum can be constructed, one need “only” solve equation (27). This, however, is still not quite the end of the story as the orbital phase $\Phi$ must be evolved simultaneously. By modelling the fluid motion in the neutron star as a set of harmonic oscillators [16, 17, 361] and incorporating mode kinetics into the Hamiltonian of the binary, [360, 225, 239] have shown how this can be achieved with a high-order PN method, though with the modes themselves calculated in full GR without Cowling (i.e. including the metric perturbations). While not all are taken into account (such as those occurring in gravities with non-metric fields where dominant mono- or dipolar radiation exists; [450]), we list PN orders at which various effects occur in Table 2. We remark in this respect that although, strictly speaking, radiation-reaction is a 2.5 PN effect, often this is said to appear at Newtonian order when using a Newtonian definition for the system’s quadrupole moment(s) (see below). To avoid confusion between different conventions for PN order counting, we focus on effects on binary motion and do not refer to the hierarchical imprints on GWs.
PN effects in shaping stellar structure have been examined in several references [e.g. 435, 388, 28] and starts already at first order, though no systematic analysis of how these directly influence the inspiral has been carried out. Tides and spins contribute hierarchically at several PN orders accounting, for instance, for the breakdown of the point-particle approximation. Tides absorb orbital energy in the Newtonian theory already (i.e. at $0^{\rm th}$ order), while spin couples to the orbital angular momentum at the 1.5th order and to self-spin and companion spin at 2nd order. At 2.5th order enters the leading-order dissipation due to GW emission [cf. equation (11)]. The electric- (magnetic-) type stationary deformations factor into the dynamics at 5th (6th) order, as introduced in Sec. IV.1.2, and the coupling between this deformation to the spin follows at the next half order. GW170817 placed constraints on each of these PN orders, as detailed in [8].
With the above in mind, the notion of dephasing can be made precise. In the stationary phase approximation, the frequency-space GW phase, $\Psi$, can be written as [e.g. 145, 101]
@@ IV.1.2 Tides: equilibrium
The equilibrium tide simply corresponds to the “$\Phi(t)=0$” portion of the dynamics. In this case, the quasiharmonic time dependence of (13) drops out, and our interest shifts to bulk, geometric deformations of the stellar surface (“zero-frequency oscillations”). The extent to which the stellar interior is suspectible to a time-independent tidal potential can be encapsulated by the (shape) Love numbers; effectively, much like the $Q_{\alpha}$ defined previously, these measure the extent of orthogonality between the stellar fluid (or solid) and some angular portion of the tidal field (i.e. $\ell,m^{\prime}$). In general, therefore, there is an infinite set of Love numbers, though often in the literature one will find the term “Love number” just to mean the quadrupolar, $m^{\prime}=0$ Love number. In the context of compact binaries in GR, Damour 103 first quantified how tidal Love numbers influence inspiral. Note also that if the star is static and axisymmetric at background order, the index $m^{\prime}$ falls out of the equations and thus the static, or even effective, Love numbers are characterised by a single harmonic number $\ell$. These reduced Love numbers are typically denoted $k_{\ell}$.
@@
which is the same $U$ from (13). Note that the multipole moments $I_{\ell m}$ are also implicitly defined at some PN order; see Section V in Thorne 415. At Newtonian order, $I_{\ell m}$ is defined through an integral over the mass-density weighted by a spherical harmonic and radius to power $\ell$:
@@
where the (axisymmetric) mass ($M_{\ell}$) and current ($S_{\ell})$ multipole moments can be defined either via the Thorne 415 or Geroch-Hansen [184] formulae. These formulations are totally equivalent by a theorem due to Gürsel 179. In the above we have the odd and even parity sectors of contractions of the trace-free Weyl tensor [e.g. 107, 310]. Using the schemes introduced by Pappas and Sotiriou 312, Suvorov and Melatos 402 to generalise the Geroch-Hansen definitions to theories beyond GR, one could try to extend the correspondence (38) to some other theory of gravity. This has not yet been attempted.
Figure 7: The main radial gravitational potential $A(R)$ — the lapse function of the effective-one-body spacetime over which the dynamics of a test particle describes the dynamics of the real binary system [70, 71] — as a function of the radial coordinate $R=a$ with $M=M_{A}+M_{B}$ (for this Figure only) in a tidal-EOB model. The binary black hole sector of this model (solid black) shows a deviation from the lapse of the Schwarzschild spacetime (dotted), indicating the dynamics of a test particle on the effective spacetime is different from that of a (plunging) orbit around a Schwarzschild black hole due to the finite symmetric mass ratio $\nu=M_{A}M_{B}/M^{2}$ of the binary (which should not be confused with the our use of $\nu$ as stellar spin used elsewhere). For the neutron star case, the gravitational potential including up to next-to-next-to-leading-order tidal effects (i.e. up to 7 PN; green dashed) is shown together with the result resumming the gravitational-self-force information up to 7.5 PN [60, red]. The specific equal-mass neutron-star binary considered here has a tidal deformability (equation 39) of $\kappa_{2}^{T}\approx 73.55$. From Bernuzzi et al. 54 with permission.
Tides effectively enhance the “attraction” between the components of the binary, leading to earlier merger relative to binary black-holes or point-particles. The tidal dynamics are elegantly described by a radial interaction potential in the effective-one-body (EOB) formalism [61, 54, 193]. PN models of tides in the EOB formalism were initially verified against numerical simulations by Baiotti et al. 43, Bernuzzi et al. 55, who demonstrated that the models can be improved with such calibrations but that it is not possible to accurately describe tides close to merger with a PN model. Such a restriction was lifted by Bernuzzi et al. 54 through resumming techniques. A schematic depiction of this attractive feature is shown in Figure 7, where the red curve describes the radial potential for a neutron-star binary with tidal resumming. A significant drop in the potential occurs at larger radius, indicating the stronger attraction felt by the binary than the equivalent binary black hole. The dominant term of the tidal imprints on waveform is delivered by the quadrupole quantity [106]
@@
for $\Lambda_{A}=\frac{2}{3}k_{2}^{A}(R_{A}/M_{A})^{5}$ with the same definition for object $B$ [see also 191, 146, 192, 105]. Note $k_{2}$ is defined in expression (34), effectively being the constant of proportionality between the quadrupole moment and the quadrupolar tidal potential (appropriately generalised to GR via the Weyl tensor; though cf. Footnote 4).